OpenStudy (anonymous):

How would I solve (x-3)(5x+4) = 0?

OpenStudy (anonymous):

Easy, one second.

OpenStudy (turingtest):

zero factor property

Parth (parthkohli):

Either (x - 3) = 0 or 5x + 4 = 0 So, x - 3 = 0 x = 3 OR 5x + 4 = 0 5x = -4 x = -4/5 Will you be able to solve the rest?

Parth (parthkohli):

(3,-4/5) is the solution for x.

OpenStudy (anonymous):

I honestly have no clue what to do. Am I just trying to plug in x?

Parth (parthkohli):

You know that 0 multiplied by any number is 0? That's what we are doing here.

Parth (parthkohli):

Essentially, we're saying that one of them(or both) has/have the value of 0.

OpenStudy (turingtest):

the zero factor property states that if we have something like\[ab=0\]then either \(a=0\) or \(b=0\) or both so we split this up and solve separately ...what parth said

OpenStudy (anonymous):

But there's only one variable which is x.. Should I just keep testing numbers until I find the one that would solve the equation to be 0??

Parth (parthkohli):

Both the solutions satisfy the equation.

OpenStudy (turingtest):

@LoyalPawn could you solve something like \[3x+5=14\]?